OpenStudy (anonymous):

what is life :( ALSO WHAT IS 43(32 X 98)= ????

OpenStudy (anonymous):

you first do what is in the parenthesis.

OpenStudy (anonymous):

32*98 = ?

OpenStudy (anonymous):

First solve 32*98 then solve 43* the sum

OpenStudy (anonymous):

So 32*98=3136 and 3136*43= 134848 so 43(32*98)= 134848

OpenStudy (anonymous):

ok